A man has twenty cons consisting of dimes and quarters. If the dimes were quarters and the quarters were dimes, he would have ninety cents more than he has now. How many dimes and quarters does he have?
Answer by stanbon(75887)

A man has twenty coins consisting of dimes and quarters. If the dimes were quarters and the quarters were dimes, he would have ninety cents more than he has now. How many dimes and quarters does he have?
--------------------------------------
EQUATIONS:
d+q = 20
Old value = New Value - 90
10d+25q = 25d+10q - 90
------------------------------
Rearrange:
d+q = 20
-15d+15q = -90
d - q = 6
----------------
2d = 26
d = 13 (# of dimes)
q = 20-d = 20-13 = 7 (# of quarters)
===========================================
Cheers,
Stan H.